OpenStudy (anonymous):

solve the inequality. -x less than or equal to 5 explain.

OpenStudy (freckles):

hint -1(-1)=1

OpenStudy (anonymous):

i just dont know how to solve it with the less than or equal sign

OpenStudy (freckles):

if you multiply both sides by a negative (or divide both sides by a negative) make sure you flip the direction of the inequality sign

OpenStudy (freckles):

just like if you have 3>2 and you multiply both sides by -1 you -3<-2

OpenStudy (anonymous):

okay so i multiply both sides by one then?

OpenStudy (freckles):

so what happens when you multiply both sides by -1?

OpenStudy (freckles):

the hint was (-1)(-1)=1 we want just x by itself 1(-1)=-1 nothing changes there so definitely multiplying both sides by -1

OpenStudy (anonymous):

5<-1?

OpenStudy (freckles):

no you have -x<=5

OpenStudy (freckles):

multiply both sides of that by -1

OpenStudy (anonymous):

oh so do i do anything else now?

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@Directrix

OpenStudy (freckles):

have you multiplied both sides by -1 yet if so can i see the result

OpenStudy (anonymous):

yeah you would get x<=-5 or 5

OpenStudy (freckles):

what is -1 times 5?

OpenStudy (anonymous):

-5

OpenStudy (freckles):

so there is no or there

OpenStudy (freckles):

what happens to the direction of the inequality

OpenStudy (anonymous):

it switches sides?

OpenStudy (freckles):

Recall if you have 3>2 and you decide you are going to multiply by -1 on both sides you have -3<-2

OpenStudy (freckles):

yes flip the inequality sign

OpenStudy (anonymous):

okay so then it would be x>-5

OpenStudy (freckles):

\[-x \le 5 \\ (-1)(-x) \ge (-1)(5) \text{ inequality sign has been flipped since we multiply/divide } \\ \text{ both sides by a negative } x \ge -5 \]

OpenStudy (anonymous):

@Sadworld

OpenStudy (freckles):

are you still stuck?

OpenStudy (freckles):

we already have the answer

OpenStudy (freckles):

x is greater than or equal to -5

OpenStudy (anonymous):

kinda like its just im not the best at math it takes me a while idk

OpenStudy (freckles):

which part confuses you

OpenStudy (anonymous):

the whole thing, its like a tiny puzzle you cant figure which goes with which

OpenStudy (freckles):

you have \[-x \le 5 \\ \text{ to get } x \text{ by itself you need } -1(-x)=x \text{ on the left } \\ \text{ \to get that you need to multiply both sides by -1 } \\ (-1)(-x) \ge (-1)(5) \\ x \ge -5 \\ \text{ nothing else to do }\]

OpenStudy (freckles):

the inequality sign changes the direction if you think back to non-variable numbers this is easy to see

OpenStudy (freckles):

you know -3<-2

OpenStudy (freckles):

if you multiply both sides by -1 you get 3>2

OpenStudy (freckles):

notice the inequality sign doesn't keep the same direction due to the multiplying/dividing by negative numbers on both sides